# Prompt to Restore, Restore to Prompt: Cyclic Prompting for Universal Adverse Weather Removal

### v1

@@ Abstract
Index Terms:
Universal adverse weather removal, Image restoration, Cyclic prompt
@@ I Introduction
Images captured in adverse weather, such as rain, snow, and fog/haze, often suffer from diminished quality, which presents significant challenges for outdoor vision systems [1, 2, 3]. Most methods primarily focus on crafting specialized networks tailored to certain weather conditions, including image deraining [4, 5, 6], defogging/dehazing [7, 8], and desnowing [9]. Some other studies [10, 11, 12] have developed a general restoration architecture and train discrete copies independently to accommodate different weather conditions. Although notable outcomes have been obtained in their intended scenarios, these task-specific models struggle to generalize effectively across different weather types and even intensity levels. This implies that tedious specialized weights are necessitated in practical applications, increasing computational costs and resources. Chen et al. [13] develop pre-trained image processing models comprising multiple heads aligned to different low-level tasks, which excavates the model capability to deal with different corruptions but requires further fine-tuning on each specific dataset.
Recent studies [14, 15, 16, 17, 18, 19] have shifted towards developing generic solutions to address multiple adverse weather challenges within a single, unified model. Although some generalizability improved on synthetic datasets, their adaptability to real-world scenarios remains limited. Inspired by the power of large-scale pre-trained vision-language models (PVLM) [20] in various computer vision tasks [21, 22], some works leverage off-the-shelf PVLMs to extract degradation-aware prompts in linguistic [23, 24] or visual [25, 26] domains and interact with the main feature flow hierarchically. For instance, DA-CLIP [23] aligns text embeddings with visual representations to predict the degradation prompt integrated into other restoration models. Liang et al. [24] construct a question prompt based on pre-trained LISA [21] to estimate the weather types and levels. Despite remarkable advancements, existing prompt-driven methods for UAWR still fall short in several limitations: 1) While PVLMs play a pivotal role in navigating weather-related content, describing weather types through simple text or converted image embeddings remains challenging, particularly in scenarios with complex weather (e.g., rain+fog, Fig. 1). 2) Visual prompts from corrupted images may not adequately express the spatial context or even introduce ambiguous semantics due to the interference of weather degradations.
@@
In summary, the main contributions of our work are
@@
  • We introduce the erase-and-paste mechanism (EPM), leveraging constrained restoration priors to embed weather-free representations into the C2P, thereby refining the restoration process. Additionally, the residue prior modulator is designed to enrich visual characteristics for better detail recovery along with EPM.
@@ II-A Task-Specific Adverse Weather Removal
Image Deraining. Rainy weather usually affects the images in two ways: rain streaks and raindrops. Fu et al. [35] propose the first convolutional neural network (CNN)-based method DerainNet for rain streak removal. As motivated, serial methods are developed that utilize advanced deep networks [36, 37, 38, 39, 40, 41, 42] or rain-related physical modeling [33, 43, 44, 45] to pursue promising results. To alleviate the limitation of CNNs in long-range dependency capturing, [5, 46, 47] combines the locality of CNNs and the global capacity of transformers, which performs favorably. Raindrops adhere to glass surfaces (e.g. windows or camera lens) and occlude the regions of background scenes. To mitigate the damage of visibility, a popular approach [48, 29, 49] is to detect the raindrop masks from rainy inputs to help recover raindrop-free ones. There are also several methods [28, 50] jointly remove the rain streaks and raindrops, significantly improving the applicability in real rainy scenes.
Image Dehazing. Hazy images often suffer from color distortion, blurry textures, and poor quality. In past years, the prevailing progress of image dehazing has been witnessed in deep learning [51, 52, 53, 54, 55, 56]. DehazeNet [51] introduces the pioneering CNN for haze removal. Considering that the haze on objects is highly related to depth, recent research [57, 58, 59] recover from the hazy image by depth estimation and depth-aware reconstruction. Tian et al. [60] integrate diffused texture priors for high-quality image restoration, which has demonstrated powerful generative ability.
Image Desnowing. Unlike rain or haze, the particle shapes and sizes in snow introduce greater complexity. From the first deep learning method DesnowNet [9], deep learning has dominated this task [30, 61, 62, 63]. DDMSNet [31] learns semantic- and geometry-aware representations utilizing semantic and depth priors. In [64], Chen et al. propose the CPLFormer that learns from cross-scale snow prototypes to handle complicated snow landscapes. Zhang et al. [65] propose the hybrid color space-based desnowing network based on the analysis of snowflake noises in the Hue channel. While the above studies acquire impressive achievements on specific weather conditions, they ignore other adverse weather which limits their scope of practical applications.
@@ III-A Overview
$\mathcal{L}=\|\tilde{\mathbf{I}}_{HQ}-\mathbf{I}_{HQ}\|_{1}+\|\hat{\mathbf{I}}_{HQ}-\mathbf{I}_{HQ}\|_{1},$ (1)
where $I_{HQ}$ denotes the ground-truth image of $\mathbf{I}_{LQ}$.
@@ III-B Composite Context Prompt
In this work, we consider the prompt in the UAWR task from two aspects to build the C2P $\mathbf{P}_{v}$. First, we integrate weather-specific knowledge $\mathbf{P}_{k}$ and learnable input-conditional vector $\mathbf{P}_{i}$ as the visual prompt $\mathbf{P}_{c}$ to perceive the weather-dedicated information. Second, the text description for the input image can be easily accessible to encapsulate the semantic information, serving as the textual prompt $\mathbf{P}_{t}$.
@@ III-C Erase-and-Paste Mechanism
Besides, in EPM, we implement residual prior extraction on the initial restored image $\tilde{\mathbf{I}}_{HQ}$ by subtracting its minimum channel from the maximum channel, i.e. residue channel, which has shown effects on rain removal [14, 42]. Such a prior $\mathbf{R}$ is weather-invariant and allows maximized maintaining the contexts. This process can be formulated by
$\mathbf{R}=\max_{c\in\{r,g,b\}}\tilde{\mathbf{I}}_{HQ}-\min_{c\in\{r,g,b\}}\tilde{\mathbf{I}}_{HQ},$ (8)
where $\max_{c\in\{r,g,b\}}$ and $\min_{c\in\{r,g,b\}}$ denotes the maximum and minimum RGB channel of $\tilde{\mathbf{I}}_{HQ}$. To take full advantage of the residual prior, in the 2nd cyclic prompt-based restoration, it complements with $\hat{\mathbf{P}}_{c}$ in the prompt block (see Fig. 3) to enrich the texture clues and structural appearance.
By combining C2P and EPM, we can form a remarkedly distinctive cyclic prompt learning-based framework, providing holistic representations such that the restoration process can be effectively adapted for adverse weather.
@@ III-D Prompt Block
$\hat{\mathbf{X}}=\mathcal{A}((\mathbf{Q},\mathbf{\mathbf{K}}),\mathbf{V})={\rm{Softmax}}(\frac{\mathbf{Q}\mathbf{K}^{\mathrm{T}}}{d})\cdot\mathbf{V},$ (9)
@@
When the current iteration is the second, the cyclic prompt $\hat{\mathbf{P}}_{c}$ by EPM and the residual prior $\mathbf{R}$ pass through the bottom branch sequentially. In addition to the cross attention between $\mathbf{X}$ and $\hat{\mathbf{P}}_{c}$ as the 1st iteration, there is a residual prior modulator (RPM) in the prompt block that conducts conditional modulation based on $\mathbf{R}$. As shown in Fig. 3(b), two $3\times 3$ convolutions followed by two cross-hierarchy information mining block (CHIMB) [83] are leveraged to learn informative residual feature $\mathbf{F}_{r}$ from $\mathbf{R}$ with large kernel attention and channel attention. Next, another two $3\times 3$ convolutional layers with two respective $1\times 1$ convolution layers are applied to generate a pair of affine transformation parameters ($\alpha$, $\beta$). Therefore, the residual prior can modulate the attentive features $\tilde{\mathbf{X}}$ as follows:
@@
$\hat{\mathbf{X}}=\begin{cases}\mathcal{A}(\mathbf{P}_{c},\mathbf{X}),&iter.=1,\\ \mathcal{R}(\mathbf{R},\mathcal{A}(\hat{\mathbf{P}}_{c},\mathbf{X}),&iter.=2,\end{cases}$ (11)
@@ IV Experiments
We conducted extensive experiments to evaluate the effectiveness of the proposed method. This section begins by outlining the experimental settings, followed by a detailed presentation of the qualitative and quantitative results. Lastly, we perform ablation studies to analyze the contributions of each key component.
@@ IV-A Implementation Details
Following the transformer setting in [10], in CyclicPrompt, from level 1 to level 4 of the encoder, the number of transformer blocks [$B_{1}$, $B_{2}$, $B_{3}$, $B_{4}$] is set to [4, 6, 6, 8] and the channel numbers are [48, 96, 192, 384]. Since the decoder employs a two-iteration design, the number of transformer blocks is reduced by half with the same channel numbers as the encoder for computational efficiency. The prompt block is incorporated before each decoder level, totaling 3 blocks in the overall CyclicPrompt network.
@@ IV-B Dataset
We conduct experiments on the All-Weather dataset [14] mainly degraded by four typical adverse weather conditions: Rain+Fog, Raindrop, and Snow. This dataset includes 18,069 images as the training data comprising 9,001 synthesized snowy images from Snow100K [9], 818 images with raindrops on windows or lenses from RainDrop [29], and 8,250 heavy rain images from Outdoor-Rain [33] for Rain+Fog. Notably, the Outdoor-Rain subset captures both rain streaks and fog-like degradation caused by rain accumulation. Due to the imbalance in the number of training samples for different weather conditions, particularly the limited size of the RainDrop subset, we perform a 10-fold resampling of the Raindrop images to ensure balanced training.

### v2

@@ I Introduction
Images captured in adverse weather, such as rain, snow, and fog/haze, often suffer from diminished quality, which presents significant challenges for outdoor vision systems [1, 2, 3]. Most methods primarily focus on crafting specialized networks tailored to certain weather conditions, including image deraining [4, 5, 6], defogging/dehazing [7, 8], and desnowing [9]. Some other studies [10, 11, 12] have developed a general restoration architecture and trained discrete copies independently to accommodate different weather conditions. Although notable outcomes have been obtained in their intended scenarios, these task-specific models struggle to generalize effectively across different weather types and even intensity levels. This implies that tedious specialized weights are necessitated in practical applications, increasing computational costs and resources. Chen et al. [13] develop pre-trained image processing models comprising multiple heads aligned to different low-level tasks, which excavates the model capability to deal with different corruptions but requires further fine-tuning on each specific dataset.
Recent studies [14, 15, 16, 17, 18, 19] have shifted towards developing generic solutions to address multiple adverse weather challenges within a single, unified model. Specifically, Li et al. [14] derive multiple encoders to learn the weather-specific representations and use the neural architecture search (NAS) technique to find the optimal one for each degradation. TransWeather [15] utilizes a shared encoder-decoder with weather-type queries to implicitly modulate degradation patterns. [17, 18] follows the multi-encoder design in [14] to accommodate diverse degradations. Although some generalizability improved on synthetic datasets, their adaptability to real-world scenarios remains limited. Inspired by the power of large-scale pre-trained vision-language models (PVLM) [20] in various computer vision tasks [21, 22], some works leverage off-the-shelf PVLMs to extract degradation-aware prompts in linguistic [23, 24] or visual [25, 26] domains and interact with the main feature flow hierarchically. For instance, DA-CLIP [23] aligns text embeddings with visual representations to predict the degradation prompt integrated into other restoration models. Liang et al. [24] construct a question prompt based on pre-trained LISA [21] to estimate the weather types and levels. Despite remarkable advancements, existing prompt-driven methods for UAWR still fall short in several limitations: 1) While PVLMs play a pivotal role in navigating weather-related content, describing weather types through simple text or converted image embeddings remains challenging, particularly in scenarios with complex weather (e.g., rain+fog, Fig. 1). 2) Visual prompts from corrupted images may not adequately express the spatial context or even introduce ambiguous semantics due to the interference of weather degradations.
@@
In summary, the main contributions of our work are:
@@
  • We introduce the erase-and-paste mechanism (EPM), leveraging constrained restoration priors to embed weather-free representations into the C2P, thereby refining the restoration process. Additionally, the residual prior modulator is designed to enrich visual characteristics for better detail recovery along with EPM.
@@ II-A Task-Specific Adverse Weather Removal
Image Deraining. Rainy weather usually affects the images in two ways: rain streaks and raindrops. Fu et al. [35] propose the first convolutional neural network (CNN)-based method DerainNet for rain streak removal. As motivated, a series of methods are developed that utilize advanced deep networks [36, 37, 38, 39, 40, 41, 42] or rain-related physical modeling [33, 43, 44, 45] to pursue promising results. For instance, [36, 37] adopt recurrent networks to progressively remove the rain by modelling the feature dependencies across multiple recurrent phases. [42] and [46] explore the residue channel prior to learn rain-free object structures for rain removal. Pan et al. propose a physics-based GAN model to ensure that the restored image is consistent with the corrupted observation. [45] investigates the physical relationship of the rain attribute and style between synthetic and real data.
To alleviate the limitation of CNNs in long-range dependency capturing, [5, 47, 48] combines the locality of CNNs and the global capacity of transformers, which performs favorably. Raindrops adhere to glass surfaces (e.g. windows or camera lens) and occlude the regions of background scenes. To mitigate the damage of visibility, a popular approach [49, 29, 50] is to detect the raindrop masks from rainy inputs to help recover raindrop-free ones. There are also several methods [28, 51] jointly remove the rain streaks and raindrops, significantly improving the applicability in real rainy scenes.
Image Dehazing. Hazy images often suffer from color distortion, blurry textures, and poor quality. In past years, the prevailing progress of image dehazing has been witnessed in deep learning [52, 53, 54, 55, 56, 57]. DehazeNet [52] introduces the pioneering CNN for haze removal. DCPDN [54] presents an edge-preserving densely connected pyramid network to estimate the transmission map in image dehazing. FFA-Net [56] devises the feature attention to facilitate the flexibility in dealing with different information. Considering that the haze on objects is highly related to depth, recent research [58, 59, 60] recover from the hazy image by depth estimation and depth-aware reconstruction. Tian et al. [61] integrate diffused texture priors for high-quality image restoration, which has demonstrated powerful generative ability.
Image Desnowing. Unlike rain or haze, the particle shapes and sizes in snow introduce greater complexity. From the first deep learning method DesnowNet [9], deep learning has dominated this task [30, 62, 63, 64]. DDMSNet [31] learns semantic- and geometry-aware representations utilizing semantic and depth priors. JSTASR [62] performs joint size and transparency-aware desnowing using partial convolutions and veiling effect removal. In [65], Chen et al. propose the CPLFormer that learns from cross-scale snow prototypes to handle complicated snow landscapes. Zhang et al. [66] propose the hybrid color space-based desnowing network based on the analysis of snowflake noises in the Hue channel. While the above studies acquire impressive achievements on specific weather conditions, they ignore other adverse weather which limits their scope of practical applications.
@@ III-A Overview
$\mathcal{L}=\|\tilde{\mathbf{I}}_{HQ}-\mathbf{I}_{HQ}\|_{1}+\lambda\|\hat{\mathbf{I}}_{HQ}-\mathbf{I}_{HQ}\|_{1},$ (1)
where $I_{HQ}$ denotes the ground-truth image of $\mathbf{I}_{LQ}$ and $\lambda$ is the hyper-parameter that controls the impact of the second iteration. We empirically set $\lambda=1$ in our experiments.
@@ III-B Composite Context Prompt
In this work, we consider the prompt in the UAWR task from two aspects to build the C2P $\mathbf{P}_{c}$. First, we integrate weather-specific knowledge $\mathbf{P}_{k}$ and learnable input-conditional vector $\mathbf{P}_{i}$ as the visual prompt $\mathbf{P}_{v}$ to perceive the weather-dedicated information. Second, the text description for the input image can be easily accessible to encapsulate the semantic information, serving as the textual prompt $\mathbf{P}_{t}$.
@@ III-C Erase-and-Paste Mechanism
Residual Channel Prior. While the initial restoration in CyclicPrompt significantly reduces degradation, several severely degraded inputs still retain residual weather artifacts in their results. To address this issue, we draw inspiration from the residual channel prior (RCP) in rain removal [14][42], which can learn weather-invariant features. In the deraining context, the colored-image intensity is expressed as
$\tilde{\mathcal{C}}(x)=\tau\rho_{w}(x)\mathbf{L}\sigma+(T-\tau)\mathbf{B}\pi$ (8)
where $T$ denotes the exposure time and the elapsed time while the weather is passing through a pixel $x$ is $\tau$. $\rho_{w}$ denotes the composition of refraction, specular reflection, and internal reflection. $\mathbf{L}=[{L_{r},L_{g},L_{b}}]^{\mathrm{T}}$ with $\mathrm{L}=L_{r}+L_{g}+L_{b}$ and $\mathbf{B}=[{B_{r},B_{g},B_{b}}]^{\mathrm{T}}$ with $\mathrm{B}=B_{r}+B_{g}+B_{b}$ denote the color vector of the light brightness and background reflection, respectively. $\sigma=\mathbf{L}/\mathrm{L}$ and $\pi=\mathbf{B}/\mathrm{B}$ represent the light chromaticity of light brightness $\mathbf{L}$ and background reflection $\mathbf{B}$, respectively. To generate the residue channel free from any degradation, one can estimate $\sigma$ via a color constancy algorithm and then apply the normalization as
$\mathcal{C}(x)=\frac{\tilde{\mathcal{C}}(x)}{\sigma}=\mathcal{C}_{w}(x)\mathbf{i}+\mathcal{C}_{bg}(x)$ (9)
where $\mathbf{i}=(1,1,1)^{\mathrm{T}}$, $\mathcal{O}_{w}=\tau\rho_{w}\mathbf{L}$, and $\mathcal{O}_{bg}=(T-\tau)\mathbf{B}/\sigma$. Notably, the light chromaticity and color effect of the spectral sensitivities are canceled during the normalization process. The achromatic nature of weather artifacts implies minimal inter-channel variance, with RGB values converging toward identity in affected regions. Therefore, the RCP can be defined by subtracting the minimum channel of the images from the maximum channel, i.e. residue channel, which can eliminate the weather influences to extract object structures. In this work, considering the foundation of typical adverse weather in the physics of atmospheric scattering, we generalize RCP to all weather types in UAWR. Therefore, we can extract the RCP $\mathbf{R}(x)$ of the initial restoration $\tilde{\mathbf{I}}_{HQ}$ by
$\mathbf{R}(x)=\tilde{\mathbf{I}}^{M}_{HQ}(x)-\tilde{\mathbf{I}}^{m}_{HQ}(x)$ (10)
where $\mathbf{I}^{M}(x)=\mathrm{max}(\mathbf{I}_{r}(x),\mathbf{I}_{g}(x),\mathbf{I}_{b}(x))$ and $\mathbf{I}^{m}(x)=\mathrm{mim}(\mathbf{I}_{r}(x),\mathbf{I}_{g}(x),\mathbf{I}_{b}(x))$. Fig. 3 shows that RCP can provide degradation-free structural contextualized information. To take full advantage of the residual prior, in the 2nd cyclic prompt-based restoration, it complements with $\hat{\mathbf{P}}_{c}$ in the prompt block (see Fig. 4) to enrich the texture clues and structural appearance. By combining C2P and EPM, we can form a distinctive cyclic prompt learning framework, providing holistic representations such that the restoration process can be effectively adapted for adverse weather.
Figure 3: Illustration of the residual prior extraction. Each row corresponds to a representative sample under different adverse weather conditions. From left to right: (1) observed LQ images; (2) the initial restoration $\tilde{\textbf{I}}_{HQ}$ by C2P; and (3) the RCP extracted from $\tilde{\textbf{I}}_{HQ}$, which preserves degradation-free structural features with enhanced edge information.
@@ III-D Prompt Block
$\hat{\mathbf{X}}=\mathcal{A}((\mathbf{Q},\mathbf{\mathbf{K}}),\mathbf{V})={\rm{Softmax}}(\frac{\mathbf{Q}\mathbf{K}^{\mathrm{T}}}{\sqrt{d}})\cdot\mathbf{V},$ (11)
@@
When the current iteration is the second, the cyclic prompt $\hat{\mathbf{P}}_{c}$ generated by EPM and the residual prior $\mathbf{R}$ pass through the bottom branch sequentially. In addition to the cross attention between $\mathbf{X}$ and $\hat{\mathbf{P}}_{c}$ as the 1st iteration, there is a residual prior modulator (RPM) in the prompt block that conducts conditional modulation based on $\mathbf{R}$. As shown in Fig. 4(b), two $3\times 3$ convolutions followed by two cross-hierarchy information mining block (CHIMB) [84] are leveraged to learn informative residual feature $\mathbf{F}_{r}$ from $\mathbf{R}$ with large kernel attention and channel attention. Next, another two $3\times 3$ convolutional layers with two respective $1\times 1$ convolution layers are applied to generate a pair of affine transformation parameters ($\alpha$, $\beta$). Therefore, the residual prior can modulate the attentive features $\tilde{\mathbf{X}}$ as follows:
@@
$\hat{\mathbf{X}}=\begin{cases}\mathcal{A}(\mathbf{P}_{c},\mathbf{X}),&iter.=1,\\ \mathcal{R}(\mathbf{R},\mathcal{A}(\hat{\mathbf{P}}_{c},\mathbf{X})),&iter.=2,\end{cases}$ (13)
@@
III-E Theoretical Analysis of Cyclic Prompt
In this work, the restoration network $f_{\theta}$ is conditioned on a prompt $\mathbf{P}$ that encodes prior knowledge. Therefore, our CyclicPrompt framework can be viewed as performing two consecutive optimizations:
$\theta_{1}$ $=\mathop{\arg\min}_{\theta}\mathcal{L}_{1}(\theta)\triangleq\|{f_{\theta}(\mathbf{I}_{LQ};\mathbf{P}_{c})-\mathbf{I}_{HQ}}\|_{1}$ (14)
$\theta_{2}$ $=\mathop{\arg\min}_{\theta}\mathcal{L}_{2}(\theta)\triangleq\|{f_{\theta}(\mathbf{I}_{LQ};\hat{\mathbf{P}}_{c},\mathbf{R})-\mathbf{I}_{HQ}}\|_{1}$
In fact, this is an iterative refinement with a better-conditioned prompt. Here, $\theta_{1}$ and $\theta_{2}$ denote the weight sets in two iterations. $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ corresponds to the two terms in Eq. (1). Intuitively, the prompt $\hat{\mathbf{P}}_{c}$ is closer to the true latent clean image manifold than $\mathbf{P}_{c}$. This is because EPM explicitly removes weather-specific information and replaces it with the weather-free textual prompt revealing the clean scene context and residual prior providing degradation-invariant structural cues (see Fig. 3). Thus, the 2nd iteration is conditioned on a more accurate prior that reduces the sensitivity to the degradation interference in LQ images, which can enhance the generalizability.
Besides, in our method, the 1st iteration can be seen to constrain the solution space $\mathcal{S}$ conditioned on the C2P $\mathbf{P}_{c}$ and the 2nd iteration with the EPM project $\mathcal{S}$ onto the manifold $\mathcal{M}$ of the clean image $\mathbf{I}_{HQ}$. Therefore, the cyclic process minimizes the manifold discrepancy $\Delta_{\mathcal{M}}$ by
$\Delta_{\mathcal{M}}=\mathrm{dist}(\mathcal{S}(\mathbf{P}_{c}),\mathcal{M})-\mathrm{dist}(\mathcal{S}(\hat{\mathbf{P}}_{c},\mathbf{R}),\mathcal{M})>0$ (15)
In this way, the C2P and EPM C2P create complementary effects and improve the optimization landscape.
@@ IV Experiments
We conducted experiments to evaluate the effectiveness of the proposed method. This section begins by outlining the experimental settings, followed by a detailed presentation of the qualitative and quantitative results. Lastly, we perform ablation studies to analyze the contributions of key components.
@@ IV-A Implementation Details
Following the transformer setting in [10], in CyclicPrompt, from level 1 to level 4 of the encoder, the number of transformer blocks $B_{1}$, $B_{2}$, $B_{3}$, $B_{4}$] is set to [4, 6, 6, 8] and the channel numbers are [48, 96, 192, 384]. Since the decoder employs a two-iteration design, the number of transformer blocks is reduced by half with the same channel numbers as the encoder for computational efficiency. The prompt block is incorporated before each decoder level, totaling 3 blocks in the overall CyclicPrompt network.
@@ IV-B Dataset
We conduct experiments on the All-Weather dataset [14] mainly degraded by four typical adverse weather conditions: Rain+Fog, Raindrop, and Snow. This dataset includes 18,069 images as the training data, comprising 9,001 synthesized snowy images from Snow100K [9], 818 images with raindrops on windows or lenses from RainDrop [29], and 8,250 heavy rain images from Outdoor-Rain [33] for Rain+Fog. Notably, the Outdoor-Rain subset captures both rain streaks and fog-like degradation caused by rain accumulation. Due to the imbalance in the number of training samples for different weather conditions, particularly the limited size of the RainDrop subset, we perform a 10-fold resampling of the Raindrop images to ensure balanced training.